OpenStudy (anonymous):

write as a single log 1/2Ln(x^2+1)-(4Ln1/2)-(1/3Ln(x-4)

OpenStudy (anonymous):

\[\frac{1}{2}\ln(x^2+1)-4\ln(\frac{1}{2})-\frac{1}{3}\ln(x-4) =\] \[\ln((x^2+1)^{\frac{1}{2}})-\ln((\frac{1}{2})^{4})-\ln((x-4)^{\frac{1}{3}}) =\] \[\ln\frac{(x^2+1)^{\frac{1}{2}}}{(\frac{1}{2})^{4}(x-4)^{\frac{1}{3}}}\] wow that took a lot longer to type out than i thought >.>

OpenStudy (anonymous):

You just need to know how to play with natural logs: \[aln(b) = \ln(a^{b}), \ln(a)+\ln(b) = \ln(a*b), \ln(a)-\ln(b) = \ln(\frac{a}{b})\]

OpenStudy (anonymous):

why is the 1/2^4 multiply with (x-4)^1/3 in the denominator

OpenStudy (anonymous):

its all subtractions isnt it?

OpenStudy (anonymous):

that is a little tricky, but i works out like this: \[\ln(a)-\ln(b)-\ln(c) = \ln(a)-(\ln(b)+\ln(c)) = \ln(a)-\ln(b*c) = \ln(\frac{a}{b*c}) \]

OpenStudy (anonymous):

i understand but dont really get it why is the negative turn into positive?

OpenStudy (anonymous):

if i had something crazy like: \[\ln(5)+\ln(10)-\ln(2)-\ln(6)+\ln(cat)-\ln(dog)\] in the end it would just be: \[\ln\frac{(5)(10)(cat)}{(2)(6)(dog)}\] Group all the positive terms and thats the numerator, then group the negative terms and thats the denominator. For the negative terms, its like you are factoring out the (-1) and adding them together, thats why you multiply them, but that (-1) is there on the outside, so you divide by that product

OpenStudy (anonymous):

hey the answer is \[Ln [16\sqrt{x^2+1}]/\sqrt[3]{x-4}\]

OpenStudy (anonymous):

where does 16 comes from?

OpenStudy (anonymous):

ah, they moved the (1/2)^4 to the top of the fraction, because: \[(\frac{1}{2})^{4} = \frac{1}{2^{4}} = \frac{1}{16}\]

OpenStudy (anonymous):

but its in the denominator, so: \[\ln\frac{(junk)}{\frac{1}{16}(otherjunk) }= \ln\frac{16(junk)}{(otherjunk)} \]

OpenStudy (anonymous):

you would multiply the top and bottom of that fraction by 16, getting rid of the fraction in the denominator, and putting a 16 in the numerator

OpenStudy (anonymous):

All-in-all this was a pretty complicated problem >.<

OpenStudy (anonymous):

yup! i understand it now! you are good thank you sir!

OpenStudy (anonymous):

how about if i have

OpenStudy (anonymous):

\[Ln^2(X-4)\]

OpenStudy (anonymous):

hmm....i dont think ive ever seen something like that....so basically its: \[\ln(\ln(x-4))\] is that correct?

OpenStudy (anonymous):

im not sure either i thought you put the 2 down = 2Ln(x-4)

OpenStudy (anonymous):

Oh, its this then: \[\ln((x-4)^{2})\] Thats when you can pull the 2 down and make it: \[2\ln(x-4)\]

OpenStudy (anonymous):

as you can see, understanding notation and interpretation is 99% of the battle lol

OpenStudy (anonymous):

hehe yea so LN^2 is same as ln((x−4)^2)

OpenStudy (anonymous):

i guess so o.O althought ive never seen: \[\ln^{2}(whatever)\] before, to each their own i suppose lol

OpenStudy (anonymous):

how about e^(-Ln3) is this = to -3?

OpenStudy (anonymous):

and e^((ln2)x) =??

OpenStudy (anonymous):

close! so close, you have the right idea, you want to cancel out the ln with e. but lets look at this term a little closer: \[-\ln(3)\] If we use that rule: \[aln(b)= \ln(b^{a})\] then we have: \[(-1)\ln(3) = \ln(3^{-1}) = \ln(\frac{1}{3})\]

OpenStudy (anonymous):

So : \[e^{-\ln(3)} = e^{\ln(\frac{1}{3})} =\frac{1}{3} \]

OpenStudy (anonymous):

hmm...for the second...lets see: \[e^{\ln(2)x} = (e^{\ln(2)})^{x} = (2)^{x} = 2^{x}\] Another way to see it is: \[e^{\ln(2)x} = e^{\ln(2^{x})} = 2^{x}\]

OpenStudy (anonymous):

is there a rule to raise the 2 to power x?

OpenStudy (anonymous):

Not really, its just an expression, or a function. not really much you can do with it, although it look a lot nicer than what we started with.

OpenStudy (anonymous):

hehe yea thanks a lot!

OpenStudy (anonymous):

no prob :)

OpenStudy (anonymous):

have a good day

OpenStudy (anonymous):

you too, good luck with your work :)

OpenStudy (anonymous):

thanks